OpenStudy (anonymous):

Let a be a given real number such as a> 2. Let the function f: R → R be defined by f (x) = | x | • (a + x) a) Find the zeropoint. Are there points where f is not differentiable? b) Determine where f increases and decrease, and provide any local and global extremal and its extreme values. c) Determine where f is convex and concave, and provide any turning points.

OpenStudy (tkhunny):

I suggest you rewrite it a little and worry about the break. \[f_{1}(x) = x\cdot(a+x)\] for x > 0 and \[f_{2}(x) = -x\cdot(a+x)\] for x < 0. I'll let you consider x = 0.

OpenStudy (anonymous):

I found zeropoits at (0,0) and (x=-a,0) , also that f is not differentiable at x=0

OpenStudy (anonymous):

the f is increasing for x>0 , but i dont how to proof how f behave for x<0

OpenStudy (tkhunny):

Did you determine the 1st Derivative away from x = 0?

OpenStudy (anonymous):

i found that f'(x)= |x|+x+a for any x but x=0 if thats right

OpenStudy (tkhunny):

Why is the absolute value still in there. For x < 0, you have a perfectly good definition WITHOUT the aboslute values.

OpenStudy (anonymous):

by using a product rule for derivatives

OpenStudy (tkhunny):

Do that for ONLY x < 0. Get rid of the absolute values.